OpenStudy (anonymous):

since f(x)=lnx at x=[1,8] satisfies the mean value theorem. How would you find c that concludes the theorem

OpenStudy (xapproachesinfinity):

you solve the equation f'(c)=f(8)-f(1)/7

OpenStudy (xapproachesinfinity):

since it satisfies the mean value thm then we conclude the above result to be true

OpenStudy (xapproachesinfinity):

the rest is easy i suppose

OpenStudy (anonymous):

so you would just set f'(c)=1/x=.3? i found the .3 in the equation you typed out

OpenStudy (xapproachesinfinity):

f'(c)=1/c the left hand side is equal to ln(8)/7 so we got 1/c=ln(8)/7 then c=7/ln(8)

OpenStudy (xapproachesinfinity):

looks like you got the right thing but you don't need to write it in decimal keep it 7/ln(8) or 7/3ln(2) unless you are asked to approximate the value in decimal places